# MetaToken: Detecting Hallucination in Image Descriptions by Meta Classification

### v1

@@ 4.5 Baseline
InstructBLIP $L$ ${92.27}^{(\pm 1.6\mathrm{e}-3)}$ $92.20^{(\pm 1.7\mathrm{e}-3)}$ ${70.68}^{(\pm 2.2\mathrm{e}-2)}$ $70.38^{(\pm 1.8\mathrm{e}-2)}$ ${21.57}^{(\pm 2.9\mathrm{e}-2)}$ $19.59^{(\pm 2.4\mathrm{e}-2)}$
$E$ ${92.30}^{(\pm 1.7\mathrm{e}-3)}$ $92.27^{(\pm 1.7\mathrm{e}-3)}$ $71.25^{(\pm 1.7\mathrm{e}-2)}$ ${{71.38}}^{(\pm 1.8\mathrm{e}-2)}$ ${23.69}^{(\pm 3.5\mathrm{e}-2)}$ $22.05^{(\pm 2.9\mathrm{e}-2)}$
Ours ${\textbf{93.08}}^{(\pm 1.2\mathrm{e}-3)}$ ${\textbf{93.08}}^{(\pm 1.3\mathrm{e}-3)}$ ${89.70}^{(\pm 4.3\mathrm{e}-3)}$ ${\textbf{89.94}}^{(\pm 6.9\mathrm{e}-3)}$ ${47.58}^{(\pm 4.7\mathrm{e}-2)}$ ${\textbf{48.14}}^{(\pm 5.0\mathrm{e}-2)}$
mPLUG-Owl $L$ ${77.64}^{(\pm 2.8\mathrm{e}-3)}$ $77.61^{(\pm 3.5\mathrm{e}-3)}$ $72.00^{(\pm 5.6\mathrm{e}-3)}$ ${72.72}^{(\pm 4.1\mathrm{e}-3)}$ ${50.87}^{(\pm 8.3\mathrm{e}-3)}$ $50.35^{(\pm 8.3\mathrm{e}-3)}$
$E$ ${78.47}^{(\pm 4.6\mathrm{e}-3)}$ ${{78.47}}^{(\pm 4.4\mathrm{e}-3)}$ $72.58^{(\pm 7.5\mathrm{e}-3)}$ ${{73.62}}^{(\pm 6.0\mathrm{e}-3)}$ ${{52.99}}^{(\pm 1.3\mathrm{e}-2)}$ ${52.55}^{(\pm 1.2\mathrm{e}-2)}$
Ours ${84.23}^{(\pm 3.5\mathrm{e}-3)}$ ${\textbf{84.89}}^{(\pm 4.3\mathrm{e}-3)}$ ${89.00}^{(\pm 3.1\mathrm{e}-3)}$ ${\textbf{90.54}}^{(\pm 2.9\mathrm{e}-3)}$ ${72.02}^{(\pm 7.8\mathrm{e}-3)}$ ${\textbf{74.78}}^{(\pm 9.0\mathrm{e}-3)}$
@@ 5.1 Hallucination Detection
In this section, we discuss the performance of our proposed method with respect to $ACC$, $AUROC$ and $AUPRC$ on four state-of-the-art LVLMs. We evaluate our method for two lightweight meta models, i.e., a logistic regression (LR) and a gradient boosting (GB) classifier. We refer to Appendix A for the configuration details. Tab. 4 summarizes our results. First, we consider the general metric set $\mathcal{M}$ (Eq. 15) for image captions based on nucleus sampling [22]. We achieve an $ACC$ of up to $91.48\%$, $AUROC$ values of up to $91.01\%$, and $AUPRC$ values of up to $78.61\%$ which clearly outperforms the one-dimensional baselines $L$ (Eq. 6) and $E$ (Eq. 10). While the gradient boosting classifier outperforms the linear model for our method in all experiments, this result does not hold for the one-dimensional baselines $L$ and $E$. Especially for the log probability baseline $L$, the linear model is superior to the GB classifier in most of the experiments. Tab. 5 states the expected calibration error ($ECE$) [48] of our method for the LR and GB classifier. Both models provide a statistical reliable classification between true and hallucinated objects.
Moreover, we evaluate our method on the extended metric set $\mathcal{M}^{\textrm{clip}}$ (Eq. 16) for the LVLMs sharing the vision encoder with CLIP [50], that is, mPLUG-Owl [67] and LLaVa [23]. The results are included in Tab. 4. The additional CLIPScore metric (Eq. 14) improves the $ACC$ by up to $2.28pp$, the $AUROC$ by up to $2.74pp$ and $AUPRC$ by up to $4.87pp$ underlining the informative content provided by the cosine similarity between the image and text embedding. However, note that the CLIPScore metric can only be calculated in a lightweight manner if the objects of interest $\mathcal{C}$ are known in advance (to calculate the text embedding $\mathcal{E}_{{Cap}_{o_{j}}}\textrm{for all possible objects}\hskip 2.84526pto_{j}\in\mathcal{C}$ beforehand) and if the CLIP image embedding $\mathcal{E}_{x}$ is calculated during the generation process. Since our focus is on a general method without any restrictions, we will stick to the more general metric set $\mathcal{M}$ (Eq. 15) in the next sections.
@@
Finally, we observe better results with respect to the $AUPRC$ for LVLMs with a higher hallucination rate, like mPLUG-Owl [67] (see Tab. 3). This behavior is expected due to the highly imbalanced datasets for strong LVLMs, that is, models with a low hallucination rate induce less positive (hallucinated) training samples which makes the problem of learning the lightweight classifier $f$ more challenging. Simultaneously, we achieve higher $ACC$ values for stronger LVLMs like InstructBLIP [8] indicating the insufficiency of the accuracy as an evaluation metric for imbalanced datasets. While, for the sake of completeness, we state the performance of our method with respect to the $ACC$, we emphasise the superior interpretability of the $AUROC$ and $AUPRC$ values for imbalanced datasets [10].
5.2 Beam Search vs. Nucleus Sampling
The results from the previous section also hold for image captions based on beam search. To compare both generation methods, we generate additional image captions using the beam search algorithm for InstructBLIP [8] and mPLUG-Owl [67]. The results are stated in Tab. 6. Again, the best results for our method are obtained by the gradient boosting classifier, clearly outperforming the single-metric baselines $L$ (Eq. 6) and $E$ (Eq. 10). As we can see from Tab. 3, beam search (beam) induces less hallucinations than nucleus sampling (sample) which is in line with previous findings [62, 34]. Thus, the resulting classification problem (beam) is more challenging, i.e. in general, we obtain better detection results for the sampling-based captions with respect to $AUROC$ and $AUPRC$ (see Tabs. 4 and 6).
While the log probability baseline $L$ (Eq. 6) shows similar classification performance for both, nucleus sampling- and beam search-based captions (see Tabs. 4 and 6), the entropy baseline $E$ (Eq. 10) is more powerful for captions generated with beam search than for sampled descriptions. Even though we observe less hallucinations using the beam search algorithm (see Tab. 3), we obtain better classification results from the entropy baseline $E$ (beam) than $E$ (sample). To this end, note that the entropy metric $E_{o_{j}}$ serves as a measure of uncertainty of generating token $t_{o_{j,s}}$ from the dictionary $\mathcal{V}$ (see Sec. 3). Thus, we expect a higher entropy, i.e., a flatter probability distribution over the dictionary for hallucinated objects and a smaller entropy for true objects. In fact, we can see in Figs. 2(c) and 2(f) that this holds for both, nucleus sampling and beam search. However, nucleus sampling induces generally smaller log probability values $L$ (sample) (see Figs. 2(a) and 2(d)), and thus, a flatter probability distribution over the dictionary than beam search resulting in a smaller variance $V$ (sample) (see Figs. 2(b) and 2(e)). Thus, for nucleus sampling, we obtain generally higher entropy values $E$ (sample) which results in a higher overlap of entropy values for true and hallucinated objects (see Fig. 2(f)), which makes the classification problem more challenging. For further insights we refer to Appendix B.
@@
Figure 2: Probability Metrics. Visualization of the probability metrics log probability $L$ (Eq. 6), variance $V$ (Eq. 9) and entropy $E$ (Eq. 10) for true and hallucinated objects. (beam) and (sample) reflect the beam search and nucleus sampling algorithm, respectively.
@@
Figure 3: LASSO Path. LASSO path for mPLUG-Owl based on nucleus sampling [22] considering the metrics set $\mathcal{M}$ (Eq. 15). $A$ denotes the maximum of the absolute values of all $G$ weight coefficients for the attention metrics $A^{g},g=0,\dots,G-1$ (Eq. 5).
5.3 Metric Analysis
@@
Table 7: Metric Rank. The average rank of the metrics $\mathcal{M}$ in the LASSO paths [12, 58] of four state-of-the art LVLMs.
avg. rank metric metric name
$1.50$ $N$ absolute occurrence (Eq. 4)
$2.00$ $A$ mean absolute attention (Eq. 5)
$3.00$ $L$ log probability (Eq. 6)
$4.00$ $C$ cumulated log probability (Eq. 7)
$5.00$ $I$ MSCOCO class index (Eq. 1)
$6.50$ $R$ variation ratio (Eq. 11)
$8.00$ $V$ variance (Eq. 9)
$8.50$ $H$ co-occurrence score (Eq. 2)
$8.75$ $E$ entropy (Eq. 10)
$9.25$ $S$ score (Eq. 8)
$10.25$ $P$ relative position (Eq. 3)
$11.50$ $M$ probability margin (Eq. 12)
$12.75$ $D$ probability difference (Eq. 13)
Figure 4: $AUPRC$ as a Function of the Number of Metrics. The classification performance of MetaToken in terms of $AUPRC$ as a function of the number of metrics for different LVLMs. The metrics are selected along the LASSO path of the respective LVLM.

### v2

@@ 4.5 Baseline
InstructBLIP $L$ $89.46^{(\pm 1.4\mathrm{e}-1)}$ $89.46^{(\pm 1.3\mathrm{e}-1)}$ $73.51^{(\pm 8.7\mathrm{e}-1)}$ $73.16^{(\pm 9.0\mathrm{e}-1)}$ $27.07^{(\pm 2.1\mathrm{e}-0)}$ $25.6^{(\pm 2.5\mathrm{e}-0)}$
$E$ $89.49^{(\pm 1.6\mathrm{e}-1)}$ $89.48^{(\pm 1.6\mathrm{e}-1)}$ $65.49^{(\pm 1.3\mathrm{e}-0)}$ $66.23^{(\pm 1.5\mathrm{e}-0)}$ $15.38^{(\pm 5.7\mathrm{e}-1)}$ $17.68^{(\pm 7.0\mathrm{e}-1)}$
Ours $91.34^{(\pm 1.7\mathrm{e}-1)}$ $\textbf{91.49}^{(\pm 1.8\mathrm{e}-1)}$ $\textbf{89.93}^{(\pm 8.9\mathrm{e}-1)}$ $\textbf{89.93}^{(\pm 7.3\mathrm{e}-1)}$ $56.07^{(\pm 1.2\mathrm{e}-0)}$ $\textbf{56.71}^{(\pm 7.6\mathrm{e}-0)}$
mPLUG-Owl $L$ $72.42^{(\pm 4.3\mathrm{e}-1)}$ $72.48^{(\pm 4.6\mathrm{e}-1)}$ $71.75^{(\pm 9.4\mathrm{e}-1)}$ $71.86^{(\pm 9.3\mathrm{e}-1)}$ $51.21^{(\pm 1.2\mathrm{e}-0)}$ $50.65^{(\pm 1.1\mathrm{e}-0)}$
$E$ $70.06^{(\pm 4.9\mathrm{e}-1)}$ $70.77^{(\pm 2.9\mathrm{e}-1)}$ $66.01^{(\pm 6.3\mathrm{e}-1)}$ $68.33^{(\pm 6.1\mathrm{e}-1)}$ $40.09^{(\pm 8.2\mathrm{e}-1)}$ $45.54^{(\pm 1.2\mathrm{e}-0)}$
Ours $82.90^{(\pm 1.9\mathrm{e}-1)}$ $\textbf{83.26}^{(\pm 2.6\mathrm{e}-1)}$ $88.41^{(\pm 3.9\mathrm{e}-1)}$ $\textbf{88.90}^{(\pm 2.8\mathrm{e}-1)}$ $75.94^{(\pm 6.2\mathrm{e}-1)}$ $\textbf{77.04}^{(\pm 5.8\mathrm{e}-1)}$
MiniGPT-4 $L$ $86.91^{(\pm 3.6\mathrm{e}-1)}$ $86.85^{(\pm 3.9\mathrm{e}-1)}$ $67.26^{(\pm 2.1\mathrm{e}-0)}$ $67.01^{(\pm 2.1\mathrm{e}-0)}$ $26.25^{(\pm 1.7\mathrm{e}-0)}$ $25.41^{(\pm 1.2\mathrm{e}-0)}$
$E$ $86.84^{(\pm 3.6\mathrm{e}-1)}$ $86.82^{(\pm 3.6\mathrm{e}-1)}$ $60.78^{(\pm 1.8\mathrm{e}-0)}$ $63.19^{(\pm 1.2\mathrm{e}-0)}$ $15.77^{(\pm 6.7\mathrm{e}-1)}$ $18.98^{(\pm 1.3\mathrm{e}-0)}$
Ours $88.92^{(\pm 3.5\mathrm{e}-1)}$ $\textbf{89.27}^{(\pm 4.9\mathrm{e}-1)}$ $88.16^{(\pm 1.5\mathrm{e}-0)}$ $\textbf{89.74}^{(\pm 1.3\mathrm{e}-0)}$ $54.90^{(\pm 6.5\mathrm{e}-0)}$ $\textbf{57.25}^{(\pm 5.7\mathrm{e}-0)}$
LLaVa $L$ $81.57^{(\pm 1.4\mathrm{e}-1)}$ $81.49^{(\pm 1.6\mathrm{e}-1)}$ $70.53^{(\pm 8.7\mathrm{e}-1)}$ $70.73^{(\pm 6.6\mathrm{e}-1)}$ $37.53^{(\pm 2.0\mathrm{e}-0)}$ $36.59^{(\pm 1.7\mathrm{e}-0)}$
$E$ $81.28^{(\pm 2.8\mathrm{e}-1)}$ $81.26^{(\pm 2.9\mathrm{e}-1)}$ $62.73^{(\pm 9.0\mathrm{e}-1)}$ $64.63^{(\pm 7.7\mathrm{e}-1)}$ $23.85^{(\pm 6.3\mathrm{e}-1)}$ $27.52^{(\pm 4.6\mathrm{e}-1)}$
Ours $87.25^{(\pm 2.0\mathrm{e}-1)}$ $\textbf{87.78}^{(\pm 3.0\mathrm{e}-1)}$ $90.05^{(\pm 4.0\mathrm{e}-1)}$ $\textbf{91.01}^{(\pm 4.3\mathrm{e}-1)}$ $70.15^{(\pm 1.0\mathrm{e}-0)}$ $\textbf{72.58}^{(\pm 1.3\mathrm{e}-0)}$
BDD100K [Yu et al., 2020]
LR GB LR GB LR GB
InstructBLIP $L$ $77.54^{(\pm 3.3\mathrm{e}-1)}$ $77.73^{(\pm 4.3\mathrm{e}-1)}$ $63.30^{(\pm 8.2\mathrm{e}-1)}$ $63.63^{(\pm 1.1\mathrm{e}-0)}$ $31.80^{(\pm 1.1\mathrm{e}-0)}$ $31.94^{(\pm 1.2\mathrm{e}-0)}$
$E$ $77.86^{(\pm 3.6\mathrm{e}-1)}$ $77.86^{(\pm 3.7\mathrm{e}-1)}$ $54.32^{(\pm 5.8\mathrm{e}-1)}$ $56.71^{(\pm 6.3\mathrm{e}-1)}$ $23.28^{(\pm 7.6\mathrm{e}-1)}$ $26.06^{(\pm 8.9\mathrm{e}-1)}$
Ours $84.07^{(\pm 2.0\mathrm{e}-1)}$ $\textbf{84.40}^{(\pm 3.1\mathrm{e}-1)}$ $87.75^{(\pm 2.8\mathrm{e}-1)}$ $\textbf{88.78}^{(\pm 3.3\mathrm{e}-1)}$ $66.04^{(\pm 1.6\mathrm{e}-0)}$ $\textbf{69.55}^{(\pm 2.0\mathrm{e}-0)}$
mPLUG-Owl $L$ $66.52^{(\pm 4.4\mathrm{e}-1)}$ $66.25^{(\pm 5.3\mathrm{e}-1)}$ $63.34^{(\pm 6.7\mathrm{e}-1)}$ $63.40^{(\pm 6.4\mathrm{e}-1)}$ $44.44^{(\pm 1.1\mathrm{e}-0)}$ $43.86^{(\pm 7.5\mathrm{e}-1)}$
$E$ $66.42^{(\pm 4.8\mathrm{e}-1)}$ $66.39^{(\pm 5.4\mathrm{e}-1)}$ $57.83^{(\pm 4.8\mathrm{e}-1)}$ $60.54^{(\pm 2.9\mathrm{e}-1)}$ $37.20^{(\pm 5.4\mathrm{e}-1)}$ $40.82^{(\pm 6.9\mathrm{e}-1)}$
Ours $78.63^{(\pm 6.4\mathrm{e}-1)}$ $\textbf{79.85}^{(\pm 6.3\mathrm{e}-1)}$ $85.06^{(\pm 6.2\mathrm{e}-1)}$ $\textbf{87.36}^{(\pm 6.7\mathrm{e}-1)}$ $69.43^{(\pm 1.6\mathrm{e}-0)}$ $\textbf{74.79}^{(\pm 2.1\mathrm{e}-0)}$
MiniGPT-4 $L$ $70.92^{(\pm 3.3\mathrm{e}-1)}$ $71.08^{(\pm 2.1\mathrm{e}-1)}$ $63.76^{(\pm 3.4\mathrm{e}-1)}$ $63.44^{(\pm 4.4\mathrm{e}-1)}$ $37.47^{(\pm 5.2\mathrm{e}-1)}$ $37.18^{(\pm 2.1\mathrm{e}-0)}$
$E$ $71.30^{(\pm 2.7\mathrm{e}-1)}$ $71.20^{(\pm 2.7\mathrm{e}-1)}$ $64.04^{(\pm 5.4\mathrm{e}-1)}$ $63.47^{(\pm 5.9\mathrm{e}-1)}$ $38.03^{(\pm 5.0\mathrm{e}-1)}$ $37.51^{(\pm 3.0\mathrm{e}-1)}$
Ours $85.83^{(\pm 2.1\mathrm{e}-1)}$ $\textbf{86.05}^{(\pm 4.9\mathrm{e}-1)}$ $90.75^{(\pm 5.1\mathrm{e}-1)}$ $\textbf{92.12}^{(\pm 4.2\mathrm{e}-1)}$ $79.57^{(\pm 1.1\mathrm{e}-0)}$ $\textbf{84.01}^{(\pm 1.3\mathrm{e}-0)}$
LLaVa $L$ $69.92^{(\pm 2.4\mathrm{e}-1)}$ $70.02^{(\pm 3.0\mathrm{e}-1)}$ $61.25^{(\pm 9.4\mathrm{e}-1)}$ $61.23^{(\pm 7.2\mathrm{e}-1)}$ $37.29^{(\pm 1.2\mathrm{e}-0)}$ $37.03^{(\pm 1.1\mathrm{e}-0)}$
$E$ $70.05^{(\pm 2.8\mathrm{e}-1)}$ $70.03^{(\pm 2.8\mathrm{e}-1)}$ $56.57^{(\pm 4.5\mathrm{e}-1)}$ $57.88^{(\pm 6.2\mathrm{e}-1)}$ $32.58^{(\pm 7.2\mathrm{e}-1)}$ $34.65^{(\pm 7.6\mathrm{e}-1)}$
Ours $80.93^{(\pm 4.6\mathrm{e}-1)}$ $\textbf{82.39}^{(\pm 3.3\mathrm{e}-1)}$ $87.46^{(\pm 3.8\mathrm{e}-1)}$ $\textbf{89.56}^{(\pm 2.1\mathrm{e}-1)}$ $70.39^{(\pm 1.5\mathrm{e}-0)}$ $\textbf{76.59}^{(\pm 8.0\mathrm{e}-1)}$
@@ 5.1 Hallucination Detection
In this section, we discuss the performance of our proposed method on four SOTA LVLMs. We evaluate our method based on an LR and GB meta model (see Sec. 8.1.2 for the configuration details). Tab. 4 summarizes the computational time for calculating the input features (Sec. 3.2), training the meta model and the prediction time for hallucination detection. As shown in Tab. 3, both models provide a calibrated classification between true and hallucinated objects reflected by a small expected calibration error ($ECE$) [Pakdaman Naeini et al., 2015]. Tab. 5 summarizes our detection results. We achieve an $ACC$ of up to $91.49\%$, $AUROC$ values of up to $92.12\%$, and $AUPRC$ values of up to $84.01\%$ which clearly outperforms the TLC-A baselines $L$ and $E$ [Petryk et al., 2023]. While the GB classifier outperforms the linear model for our method in all experiments, this result does not hold for the one-dimensional baselines $L$ and $E$. Especially for the log probability-based token-level confidence $L$, the linear model is superior to the GB classifier in most of the experiments.
@@
Moreover, we observe better detection results with respect to $AUPRC$ on the BDD100K dataset than on the MSCOCO data. This behavior is expected since the MSCOCO dataset is widely used as an instruction tuning dataset for pre-trained LVLMs leading to lower hallucination rates on MSCOCO (see Tab. 2). Thus, the LVLMs induce less positive (hallucinated) training samples when generating image captions, which makes the problem of learning the lightweight classifier $f$ more challenging. Simultaneously, we achieve higher $ACC$ values on the MSCOCO dataset indicating the insufficiency of the $ACC$ as an evaluation metric for imbalanced datasets. While, for the sake of completeness, we state the performance of our method with respect to the $ACC$, we emphasize the superior interpretability of the $AUROC$ and $AUPRC$ values for imbalanced datasets [Davis and Goadrich, 2006].
Table 6: Experimental Results based on Beam Search [Bisiani, 1987]. Hallucination detection results on SOTA LVLMs. Ours refers to the feature set $\mathcal{M}$. The best results in each block are highlighted.
MSCOCO [Lin et al., 2014]
LR GB LR GB LR GB
InstructBLIP $L$ $92.27^{(\pm 1.6\mathrm{e}-1)}$ $92.19^{(\pm 1.7\mathrm{e}-1)}$ $70.68^{(\pm 2.2e+0)}$ $70.37^{(\pm 1.8e+0)}$ $21.57^{(\pm 2.9e+0)}$ $19.56^{(\pm 2.5e+0)}$
$E$ $92.30^{(\pm 1.7\mathrm{e}-1)}$ $92.27^{(\pm 1.7\mathrm{e}-1)}$ $71.25^{(\pm 1.7e+0)}$ $71.38^{(\pm 1.8e+0)}$ $23.69^{(\pm 3.5e+0)}$ $22.05^{(\pm 2.9e+0)}$
Ours $\textbf{93.07}^{(\pm 1.3\mathrm{e}-1)}$ $93.05^{(\pm 1.4\mathrm{e}-1)}$ $89.69^{(\pm 3.8\mathrm{e}-1)}$ $\textbf{89.74}^{(\pm 5.9\mathrm{e}-1)}$ $\textbf{47.45}^{(\pm 4.9e+0)}$ $47.40^{(\pm 5.0e+0)}$
mPLUG-Owl $L$ $77.64^{(\pm 2.8\mathrm{e}-1)}$ $77.61^{(\pm 3.5\mathrm{e}-1)}$ $72.00^{(\pm 5.6\mathrm{e}-1)}$ $72.71^{(\pm 4.0\mathrm{e}-1)}$ $50.87^{(\pm 8.3\mathrm{e}-1)}$ $50.35^{(\pm 8.4\mathrm{e}-1)}$
$E$ $78.47^{(\pm 4.6\mathrm{e}-1)}$ $78.47^{(\pm 4.4\mathrm{e}-1)}$ $72.58^{(\pm 7.5\mathrm{e}-1)}$ $73.62^{(\pm 6.0\mathrm{e}-1)}$ $52.99^{(\pm 1.3e+0)}$ $52.55^{(\pm 1.2e+0)}$
Ours $84.10^{(\pm 3.4\mathrm{e}-1)}$ $\textbf{84.57}^{(\pm 4.5\mathrm{e}-1)}$ $88.85^{(\pm 3.4\mathrm{e}-1)}$ $\textbf{89.54}^{(\pm 3.7\mathrm{e}-1)}$ $71.89^{(\pm 9.6\mathrm{e}-1)}$ $\textbf{73.03}^{(\pm 1.1e+0)}$
@@
Figure 2: Probability Features. Visualization of the probability features log probability $L$ (Eq. 4), variance $V$ (Eq. 7) and entropy $E$ (Eq. 8) for true and hallucinated objects. (beam) and (sample) reflect the beam search and nucleus sampling algorithm, respectively.
5.2 Beam Search vs. Nucleus Sampling
The results from the previous section also hold for image captions based on beam search. To compare both generation methods, we generate additional image captions for MSCOCO using the beam search algorithm for InstructBLIP [Dai et al., 2023a] and mPLUG-Owl [Ye et al., 2023]. The results are stated in Tab. 6. Again, generally the best results for our method are obtained by the gradient boosting classifier, clearly outperforming the single-feature baselines $L$ (Eq. 4) and $E$ (Eq. 8). As we can see from Tab. 2, beam search (beam) induces less hallucinations than nucleus sampling (sample) which is in line with previous findings [Wang et al., 2023, Li et al., 2022]. Thus, the resulting classification problem (beam) is more challenging, i.e., in general we obtain better detection results for the sampling-based captions with respect to $AUROC$ and $AUPRC$ (see Tabs. 5 and 6).
While the log probability baseline $L$ (Eq. 4) shows similar classification performance for both, nucleus sampling- and beam search-based captions (see Tabs. 5 and 6), the entropy baseline $E$ (Eq. 8) is more powerful for captions generated with beam search than for sampled descriptions. Even though we observe less hallucinations using the beam search algorithm (see Tab. 2), we obtain better classification results from the entropy baseline $E$ (beam) than $E$ (sample). To this end, note that the entropy feature $E_{o_{j}}$ serves as a measure of uncertainty of generating token $t_{o_{j,s}}$ from the dictionary $\mathcal{V}$ (see Sec. 3). Thus, we expect a higher entropy, i.e., a flatter probability distribution over the dictionary for hallucinated objects and a smaller entropy for true objects. In fact, we can see in Figs. 2(c) and 2(f) that this holds for both, nucleus sampling and beam search. However, nucleus sampling induces generally smaller log probability values $L$ (sample) (see Figs. 2(a) and 2(d)), and thus, a flatter probability distribution over the dictionary than beam search resulting in a smaller variance $V$ (sample) (see Figs. 2(b) and 2(e)). Thus, for nucleus sampling, we obtain generally higher entropy values $E$ (sample) which results in a higher overlap of entropy values for true and hallucinated objects (see Fig. 2(f)), which makes the classification problem more challenging. For further insights we refer to Sec. 8.4.
@@
Figure 3: LASSO Path. LASSO path for $\mathcal{M}$. $A$ denotes the maximum of the absolute values of all $G$ weight coefficients for the attention features $A^{g},g=0,\dots,G-1$.
@@
Table 7: Feature Rank. The average rank of the features $\mathcal{M}$ in the LASSO paths of four SOTA art LVLMs.
avg. rank feature feature name
$1.375$ $N$ absolute occurrence (Eq. 2)
$2.125$ $A$ mean absolute attention (Eq. 3)
$3.625$ $C$ cumulated log probability (Eq. 5)
$5.125$ $L$ log probability (Eq. 4)
$6.875$ $R$ variation ratio (Eq. 9)
$6.875$ $V$ variance (Eq. 7)
$7.250$ $E$ entropy (Eq. 8)
$7.625$ $D$ probability difference (Eq. 11)
$8.000$ $S$ score (Eq. 6)
$8.375$ $M$ probability margin (Eq. 10)
$8.750$ $P$ relative position (Eq. 1)
5.3 Feature Analysis